# Mapping Degeneration Meets Label Evolution: Learning Infrared Small Target Detection with Single Point Supervision

### v2

@@ 5.1 Implementation Details
Three public datasets NUAA-SIRST [8], NUDT-SIRST [24], and IRSTD-1K [54] are used in our experiments. We followed [24] to split the training and test sets of NUAA-SIRST and NUDT-SIRST, and followed [54] to split IRSTD-1K. We employed two pixel-level metrics (i.e., intersection over union ($IoU$) and pixel accuracy ($PA$)) and two target-level metrics (i.e., probability of detection ($P_{d}$) and false-alarm rate ($F_{a}$)) for performance evaluation.
@@ 5.2.1 Analyses of Mapping Degeneration
Table 1: Average $IoU$ (×$\times 10^{2}$), $P_{d}$ (×$\times 10^{2}$) and $F_{a}$(×$\times 10^{6}$) values on NUAA-SIRST[8] NUDT-SIRST [24] and IRSTD-1K [54] achieved by DNA-Net with (w/) and without (w/o) LESPS under centroid, coarse point supervision together with full supervision.
@@ 5.2.2 Analyses of the Label Evolution Framework
Effectiveness. We compare the average results of NUAA-SIRST[8], NUDT-SIRST [24], and IRSTD-1K [54] datasets achieved by DNA-Net with (i.e., w/) and without (i.e., w/o) LESPS under centroid and coarse point supervision, respectively. Note that, the results of DNA-Net w/o LESPS are calculated under extremely low threshold [^6] With point supervision, the results of DNA-Net w/o LESPS calculated under threshold 0.5 are all zeros. (i.e., 0.15) while those of DNA-Net w/ LESPS are calculated under the standard threshold (i.e., 0.5 [24, 54]). As shown in Table 1, as compared to full supervision, the results of DNA-Net w/o LESPS are extremely low, which demonstrates that SIRST detection on single point supervision is a challenging task. In contrast, DNA-Net w/ LESPS can achieve significant performance improvement under both point supervisions in terms of all the metrics, which approximate the performance of their fully supervised counterparts. Note that, $P_{d}$ results of DNA-Net w/o LESPS under coarse point supervision are over half lower than those under the centroid ones, while the results of DNA-Net w/ LESPS under these two kinds of point supervision are comparable. It demonstrates that LESPS can generalize well to manual annotation errors.
@@
Table 2: $IoU$ (×$\times 10^{2}$), $P_{d}$ (×$\times 10^{2}$) and $F_{a}$(×$\times 10^{6}$) values of different methods achieved on NUAA-SIRST[8], NUDT-SIRST [24] and IRSTD-1K [54] datasets. “CNN Full”, “CNN Centroid”, and “CNN Coarse” represent CNN-based methods under full supervision, centroid and coarse point supervision. “+” represents CNN-based methods equipped with LESPS.
Methods Description NUAA-SIRST[8] NUDT-SIRST [24] IRSTD-1K [54] Average
@@
Figure 9: Visual detection results of different methods achieved on NUAA-SIRST[8], NUDT-SIRST [24] and IRSTD-1K [54] datasets. Correctly detected targets and false alarms are highlighted by red and orange circles, respectively.
@@ 5.3 Comparison to State-of-the-art Methods
Comparison to SISRT detection methods. We apply our LESPS to several state-of-the-art CNN-based methods, including ACM [7], ALCNet [8] and DNA-Net [24]. For fair comparisons, we retrained all models on NUAA-SIRST [8], NUDT-SIRST [24], and IRSTD-1K [54] datasets with the same settings. In addition, we add the results of six fully supervised CNN-based methods (MDvsFA [45], ACM [7], ALCNet [8], DNA-Net [24], ISNet [54], UIU-Net [46]) and six traditional methods (Top-Hat [40], RLCM [15], TLLCM [16], MSPCM [17], IPI [11], PSTNN [51]) as the baseline results.
@@
Table 3: Average $IoU$ (×$\times 10^{2}$), $P_{d}$ (×$\times 10^{2}$), $F_{a}$(×$\times 10^{6}$) values on NUAA-SIRST[8], NUDT-SIRST [24] and IRSTD-1K [54] datasets of DNA-Net trained with pseudo labels generated by input intensity threshold, LCM-based methods [15, 34, 16] and LESPS under centroid and coarse point supervision.
@@ 5.4 Discussion of Loss Function
Table 4: Average $IoU$(×$\times 10^{2}$), $P_{d}$(×$\times 10^{2}$), $F_{a}$(×$\times 10^{3}$) values on NUAA-SIRST[8], NUDT-SIRST [24] and IRSTD-1K [54] datasets of different methods when computing negative loss on $i$ handcrafted (handi), $j$ randomly sampled (randj) and all background points.
@@ B Visual Results of Labels and Network Predictions
In this section, we provides additional visual results of regressed labels during training and network predictions during inference on NUAA-SIRST[8], NUDT-SIRST [24], and IRSTD-1K [54] datasets. It can be observed from Figure II that our LESPS can effectively regress mask labels during training, and can achieve accurate pixel-level SIRST detection in the inference stage.
@@ D.1 Comparison to SISRT Detection Methods
Table III: $IoU$ (×$\times 10^{2}$), $P_{d}$ (×$\times 10^{2}$) and $F_{a}$(×$\times 10^{6}$) values of different methods achieved on NUAA-SIRST[8] NUDT-SIRST [24] and IRSTD-1K [54] datasets. “CNN Full”, “CNN Centroid”, and “CNN Coarse” represent CNN-based methods under full supervision, centroid and coarse point supervision. “+” represents CNN-based methods equipped with LESPS.
Methods Description NUAA-SIRST[8] NUDT-SIRST [24] IRSTD-1K [54] Average
@@
Figure IV: ROC results of different methods achieved on (a) NUAA-SIRST[8], (b) NUDT-SIRST [24], and (c) IRSTD-1K [54] datasets. “Centroid+” represents CNN-based methods equipped with LESPS under centroid point supervision.
Figure V: Visual detection results of different methods achieved on NUAA-SIRST[8], NUDT-SIRST [24], and IRSTD-1K [54] datasets. Correctly detected targets and false alarms are highlighted by red and orange circles, respectively.
Figure IV provides ROC results of ACM [7], ALCNet [8], DNA-Net [24] equipped with LESPS under centroid point supervision (i.e., ACM Centroid+, ALCNet Centroid+, DNA-Net Centroid+) and their fully supervised counterparts (i.e., ACM, ALCNet, DNA-Net) achieved on NUAA-SIRST[8], NUDT-SIRST [24], and IRSTD-1K [54] datasets. It can be observed that ROC results of ACM Centroid+, ALCNet Centroid+, DNA-Net Centroid+, and ACM, ALCNet, DNA-Net only have minor differences (i.e., less than 5%).

### v3

@@ 5.1 Implementation Details
Three public datasets SIRST [8], NUDT-SIRST [24], and IRSTD-1K [54] are used in our experiments. We followed [24] to split the training and test sets of SIRST and NUDT-SIRST, and followed [54] to split IRSTD-1K. We employed two pixel-level metrics (i.e., intersection over union ($IoU$) and pixel accuracy ($PA$)) and two target-level metrics (i.e., probability of detection ($P_{d}$) and false-alarm rate ($F_{a}$)) for performance evaluation.
@@ 5.2.1 Analyses of Mapping Degeneration
Table 1: Average $IoU$ (×$\times 10^{2}$), $P_{d}$ (×$\times 10^{2}$) and $F_{a}$(×$\times 10^{6}$) values on SIRST[8] NUDT-SIRST [24] and IRSTD-1K [54] achieved by DNA-Net with (w/) and without (w/o) LESPS under centroid, coarse point supervision together with full supervision.
@@ 5.2.2 Analyses of the Label Evolution Framework
Effectiveness. We compare the average results of SIRST[8], NUDT-SIRST [24], and IRSTD-1K [54] datasets achieved by DNA-Net with (i.e., w/) and without (i.e., w/o) LESPS under centroid and coarse point supervision, respectively. Note that, the results of DNA-Net w/o LESPS are calculated under extremely low threshold [^6] With point supervision, the results of DNA-Net w/o LESPS calculated under threshold 0.5 are all zeros. (i.e., 0.15) while those of DNA-Net w/ LESPS are calculated under the standard threshold (i.e., 0.5 [24, 54]). As shown in Table 1, as compared to full supervision, the results of DNA-Net w/o LESPS are extremely low, which demonstrates that SIRST detection on single point supervision is a challenging task. In contrast, DNA-Net w/ LESPS can achieve significant performance improvement under both point supervisions in terms of all the metrics, which approximate the performance of their fully supervised counterparts. Note that, $P_{d}$ results of DNA-Net w/o LESPS under coarse point supervision are over half lower than those under the centroid ones, while the results of DNA-Net w/ LESPS under these two kinds of point supervision are comparable. It demonstrates that LESPS can generalize well to manual annotation errors.
@@
Table 2: $IoU$ (×$\times 10^{2}$), $P_{d}$ (×$\times 10^{2}$) and $F_{a}$(×$\times 10^{6}$) values of different methods achieved on SIRST[8], NUDT-SIRST [24] and IRSTD-1K [54] datasets. “CNN Full”, “CNN Centroid”, and “CNN Coarse” represent CNN-based methods under full supervision, centroid and coarse point supervision. “+” represents CNN-based methods equipped with LESPS.
Methods Description SIRST[8] NUDT-SIRST [24] IRSTD-1K [54] Average
@@
Figure 9: Visual detection results of different methods achieved on SIRST[8], NUDT-SIRST [24] and IRSTD-1K [54] datasets. Correctly detected targets and false alarms are highlighted by red and orange circles, respectively.
@@ 5.3 Comparison to State-of-the-art Methods
Comparison to SISRT detection methods. We apply our LESPS to several state-of-the-art CNN-based methods, including ACM [7], ALCNet [8] and DNA-Net [24]. For fair comparisons, we retrained all models on SIRST [8], NUDT-SIRST [24], and IRSTD-1K [54] datasets with the same settings. In addition, we add the results of six fully supervised CNN-based methods (MDvsFA [45], ACM [7], ALCNet [8], DNA-Net [24], ISNet [54], UIU-Net [46]) and six traditional methods (Top-Hat [40], RLCM [15], TLLCM [16], MSPCM [17], IPI [11], PSTNN [51]) as the baseline results.
@@
Table 3: Average $IoU$ (×$\times 10^{2}$), $P_{d}$ (×$\times 10^{2}$), $F_{a}$(×$\times 10^{6}$) values on SIRST[8], NUDT-SIRST [24] and IRSTD-1K [54] datasets of DNA-Net trained with pseudo labels generated by input intensity threshold, LCM-based methods [15, 34, 16] and LESPS under centroid and coarse point supervision.
@@ 5.4 Discussion of Loss Function
Table 4: Average $IoU$(×$\times 10^{2}$), $P_{d}$(×$\times 10^{2}$), $F_{a}$(×$\times 10^{3}$) values on SIRST[8], NUDT-SIRST [24] and IRSTD-1K [54] datasets of different methods when computing negative loss on $i$ handcrafted (handi), $j$ randomly sampled (randj) and all background points.
@@ B Visual Results of Labels and Network Predictions
In this section, we provides additional visual results of regressed labels during training and network predictions during inference on SIRST[8], NUDT-SIRST [24], and IRSTD-1K [54] datasets. It can be observed from Figure II that our LESPS can effectively regress mask labels during training, and can achieve accurate pixel-level SIRST detection in the inference stage.
@@ D.1 Comparison to SISRT Detection Methods
Table III: $IoU$ (×$\times 10^{2}$), $P_{d}$ (×$\times 10^{2}$) and $F_{a}$(×$\times 10^{6}$) values of different methods achieved on SIRST[8] NUDT-SIRST [24] and IRSTD-1K [54] datasets. “CNN Full”, “CNN Centroid”, and “CNN Coarse” represent CNN-based methods under full supervision, centroid and coarse point supervision. “+” represents CNN-based methods equipped with LESPS.
Methods Description SIRST[8] NUDT-SIRST [24] IRSTD-1K [54] Average
@@
Figure IV: ROC results of different methods achieved on (a) SIRST[8], (b) NUDT-SIRST [24], and (c) IRSTD-1K [54] datasets. “Centroid+” represents CNN-based methods equipped with LESPS under centroid point supervision.
Figure V: Visual detection results of different methods achieved on SIRST[8], NUDT-SIRST [24], and IRSTD-1K [54] datasets. Correctly detected targets and false alarms are highlighted by red and orange circles, respectively.
Figure IV provides ROC results of ACM [7], ALCNet [8], DNA-Net [24] equipped with LESPS under centroid point supervision (i.e., ACM Centroid+, ALCNet Centroid+, DNA-Net Centroid+) and their fully supervised counterparts (i.e., ACM, ALCNet, DNA-Net) achieved on SIRST [8], NUDT-SIRST [24], and IRSTD-1K [54] datasets. It can be observed that ROC results of ACM Centroid+, ALCNet Centroid+, DNA-Net Centroid+, and ACM, ALCNet, DNA-Net only have minor differences (i.e., less than 5%).